There is a certain chemical reaction that involves two reactants producing one product. A perfect chemist performs this reaction using 50.0 grams of each reactant, and produced 82.0 grams of product.
What is the amount, in grams, of the excess reactant?
1
Expert's answer
2020-06-15T14:31:15-0400

This reaction might be described by the following equation:

A + B = C

Every chemical equation adheres to the law of conservation of mass, which states that m (products) = m (reactants)

We've got 50.0 + 50.0 = 100.0 g of reactants

They produce only 82.0 g of products

This means that mass of chemicals reacted in fact was also 82.0 g

100.0 - 82.0 = 18.0 g is excess.